A study compared five different methods for teaching descriptive statistics. The five methods were traditional lecture and discussion, programmed textbook instruction, programmed text with lectures, computer instruction, and computer instruction with lectures. 45 students were randomly assigned, 9 to each method. After completing the course, students took a 1-hour exam.

Which of the following is the correct degrees of freedom for an F-test for evaluating if the average test scores are different for the different teaching methods?

A. dfG = 40, dfE = 4
B. dfG = 4, dfE = 40
C. dfG = 5, dfE = 45
D. dfG = 4, dfE = 44
E. dfG = 45, dfE = 4
1
Expert's answer
2020-06-10T17:33:29-0400

"Given\; that, n=45,k=5,then,\\
 dfG = k-1=5-1=4\\
 dfE = n-k=45-5=40\\
\therefore\text{The correct answer is:}\\
B) dfG = 4, dfE = 40"